# Short-Context dominance: How Much Local Context Natural Language Actually Needs?

### v1

@@ D.1 Motivations on DaMCL
    Using the Top-1 token from the sampling distribution is not always a reliable way to evaluate next-token prediction, as greedy decoding often results in low-quality or repetitive outputs (holtzman2020curious). More recent sampling strategies instead aim to identify a set of valid next tokens (zhu2024adaptive; zhou2025balancingdiversityriskl), shifting the focus away from single-token probabilities toward broader support coverage.
@@ D.2 Additional Metrics for DaMCL
These metrics are widely employed in applications such as knowledge distillation and in assessing performance degradation of large language models under various conditions (miniLLM_KD; tailoringLLMGen; advmomentDistillation).
@@ E.1.1 Controlled Validation Experiment on LongEval Task
Using the LongEval benchmark (longchat2023), we construct prompts containing multiple register lines, each with a line_id and <REGISTER_CONTENT> value, where the model must identify specific content. We create two types of queries with known ground-truth labels: short-context queries where the answer (a line_id) appears in the last 32 tokens, and long-context queries where the answer (a <REGISTER_CONTENT>) requires information from the full context. We conduct the experiments on three models: LLaMA-3-8B (grattafiori2024llama3herdmodels), Mistral-7B-Instruct (v0.2) (jiang2023mistral7b), and Qwen2-7B (yang2024qwen2technicalreport). As shown in Figure 23, across all models, we observe the same distributional trend, confirming the robustness of the results.
@@ E.1.2 Needle-in-a-Haystack Experiment on General Text
We next adapt a needle-in-a-haystack (NIAH) style setup to general text. Following the classic needle-in-a-haystack test by KamradtNIAH2023, given a natural context, we insert a needle statement with a randomly generated 6-digit number: "The magic number is xxxxxx" at a specific position in the text. At the end of the document, we append a query prompt: "The magic number mentioned in the provided text is ", expecting the model to output the correct number from the needle. If the needle statement is placed within the final 32 tokens, the answer is recoverable from the local suffix, and we label the case as a short context. Otherwise, when the needle is inserted far away, answering requires long-range recall, and we label it as a long context. We only retain cases where the model outputs the correct number, ensuring the prediction truly relies on the inserted statement rather than hallucination.
@@ E.1.3 Additional Long-short context detection on LongEval
In addition to the setup described above, which for the most part follows a similar structure to that of (PerplexityLongCtx), we consider a different setup where instead of using the line_id as the token representing a short context token, we use <REGISTER_CONTENT> in a recent line (less than 32 tokens away). Compared to above, the long-context queries remain the same, but for the short-context queries we use a <REGISTER_CONTENT> from a recent line. This allows us to have a more direct comparison between long and short context as the nature of both tokens are the same (same style of registered content as opposed to comparing to line-id) while the positions of the reference line in the local or long context position changes their LSDS behavior.
@@ E.1.4 Prior-Work Oracle (LSD/LCL) on General Text
As described in Section 5, we evaluate a prior-work oracle detector that follows the log-probability intervention from PerplexityLongCtx. The oracle Long-Short Difference (LSD) and Long-Context Likelihood (LCL) are defined as follows:
@@ E.2 Ablation on Nucleus Sampling Parameter ($p$)
In this section, we examine how varying the nucleus sampling parameter $p$ influence our LSDS and the resulting JSD threshold. We report results for 4 settings: no nucleus sampling, $p=0.95$, $p=0.90$, and $p=0.80$. We conduct the experiment with Mistral-7B-Instruct (v0.2) (jiang2023mistral7b) on 5000 contexts randomly selected from Government Report (huang2021efficient), with full context length ranging from 100–1000 tokens. The ground truth oracle we use here is Prior Work Oracle as introduced in Sec. 5.
@@
To quantitatively assess performance, we compute the maximized Youden’s $J$ statistic (youden1950index):
@@
The setting $p=0.90$ achieves the highest Youden’s $J$ ($0.840$), indicating the most balanced trade-off between sensitivity (TPR) and specificity ($1-\text{FPR}$). While $p=0.80$ achieves the highest accuracy (91.9%), its $J$ value is lower, reflecting weaker overall discriminative balance. Therefore, we adopt $p=0.90$ as our standard nucleus sampling configuration in the main experiments. This choice is consistent with prior work recommending $p$ in the 0.9–0.95 range for stable yet diverse generation quality (holtzman2020curious).
@@ F.1 Probability Shift as Selection Metric
A comparison between this form of metric and our probability difference LSPS becomes important when you consider works such as PerplexityLongCtx use the log probability ratio to detect long-short context/next-tokens. Additionally, in CoherenceBoosting and CADBoost the authors use the log ratio for probability and logits with in the form of $\mathbf{p}(\mathbf{s})_{t}^{1.5}\mathbf{p}_{[-32:]}(\mathbf{s})_{t}^{0.5}(t)$ to improve the final probability of long-context tokens. Implicitly, such methods select and increase the probability of tokens who’s probability ratio increases rather than the difference itself.
In order to compare the choice of metric, we compare the use of LSPS and LSPR when boosting the final word/token of on the validation set of LAMBADA (lambada) dataset. For each context, we have a ground truth next token, which we call the target token $\hat{t}$. All other tokens we call incorrect tokens. We analyze the choice of which token’s are selected for probability increase rather than the improvement itself, therefore we don’t include multiplicative factors like $\lambda$. For each case of boosting, we consider the following scenarios:
@@
On the other hand, for the probability difference case, we have implicitly set a minimal lower bound on the $\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s})_{t}\geq\epsilon$ by only selecting tokens where $|\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s})_{t}-\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s}_{[-32:]})_{t}|\geq\epsilon$. This allows us to circumvent tokens who have very low probability where the log ratio behavior could be noisy and misleading. We believe this is a potential reason for why the long context token selection algorithm in PerplexityLongCtx uses $\log\left[\mathbf{p}(\mathbf{s})\right]_{t}\geq 10^{-2}$ ( Long-Context Likelihood (LCL) ) in addition to their log probability ratio to classify tokens.
@@ F.4 Impact of Decoding
F.5 Analysis on recent datasets/models
We provide additional MCL results on more recent models, including Qwen-3-8B Base (yang2025qwen3technicalreport) and Mistral-v0.3-7B (mistralai_mistral7bv03), evaluated on both short- and long-context datasets. These results, shown in Fig. 34, confirm that short-context dominance persists even in these updated model families. We further evaluate TaBoo on two additional QA datasets from L-Eval (an2023leval), namely NaturalQuestions, MultiDocQA and ScientificQA, using the same models. To ensure that the QA tasks genuinely require contextual information (i.e., cannot be answered in a closed-book setting), we first run each model on the question alone and remove any example for which the model achieves an F1 score of $\geq 0.4$. As shown in Tab. 10, both models find these QA tasks challenging due to their longer contexts and higher difficulty. Nonetheless, TaBoo consistently improves performance over vanilla inference and CAD, demonstrating that our boosting approach remains effective even on these stronger, more context-heavy benchmarks.

### v2

@@ D.1 Motivations on DaMCL
    Using the Top-1 token from the sampling distribution is not always a reliable way to evaluate next-token prediction, as greedy decoding often results in low-quality or repetitive outputs (Holtzman et al., 2020). More recent sampling strategies instead aim to identify a set of valid next tokens (Zhu et al., 2024; Zhou et al., 2025), shifting the focus away from single-token probabilities toward broader support coverage.
@@ D.2 Additional Metrics for DaMCL
These metrics are widely employed in applications such as knowledge distillation and in assessing performance degradation of large language models under various conditions (Gu et al., 2024; Ji et al., 2023a; Jia, 2024).
@@ E.1.1 Controlled Validation Experiment on LongEval Task
Using the LongEval benchmark (Li* et al., 2023), we construct prompts containing multiple register lines, each with a line_id and <REGISTER_CONTENT> value, where the model must identify specific content. We create two types of queries with known ground-truth labels: short-context queries where the answer (a line_id) appears in the last 32 tokens, and long-context queries where the answer (a <REGISTER_CONTENT>) requires information from the full context. We conduct the experiments on three models: LLaMA-3-8B (Grattafiori et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023), and Qwen2-7B (Yang et al., 2024). As shown in Figure 23, across all models, we observe the same distributional trend, confirming the robustness of the results.
@@ E.1.2 Needle-in-a-Haystack Experiment on General Text
We next adapt a needle-in-a-haystack (NIAH) style setup to general text. Following the classic needle-in-a-haystack test by Kamradt (2023), given a natural context, we insert a needle statement with a randomly generated 6-digit number: "The magic number is xxxxxx" at a specific position in the text. At the end of the document, we append a query prompt: "The magic number mentioned in the provided text is ", expecting the model to output the correct number from the needle. If the needle statement is placed within the final 32 tokens, the answer is recoverable from the local suffix, and we label the case as a short context. Otherwise, when the needle is inserted far away, answering requires long-range recall, and we label it as a long context. We only retain cases where the model outputs the correct number, ensuring the prediction truly relies on the inserted statement rather than hallucination.
@@ E.1.3 Additional Long-short context detection on LongEval
In addition to the setup described above, which for the most part follows a similar structure to that of (Fang et al., 2025), we consider a different setup where instead of using the line_id as the token representing a short context token, we use <REGISTER_CONTENT> in a recent line (less than 32 tokens away). Compared to above, the long-context queries remain the same, but for the short-context queries we use a <REGISTER_CONTENT> from a recent line. This allows us to have a more direct comparison between long and short context as the nature of both tokens are the same (same style of registered content as opposed to comparing to line-id) while the positions of the reference line in the local or long context position changes their LSDS behavior.
@@ E.1.4 Prior-Work Oracle (LSD/LCL) on General Text
As described in Section 5, we evaluate a prior-work oracle detector that follows the log-probability intervention from Fang et al. (2025). The oracle Long-Short Difference (LSD) and Long-Context Likelihood (LCL) are defined as follows:
@@ E.2 Ablation on Nucleus Sampling Parameter ($p$)
In this section, we examine how varying the nucleus sampling parameter $p$ influence our LSDS and the resulting JSD threshold. We report results for 4 settings: no nucleus sampling, $p=0.95$, $p=0.90$, and $p=0.80$. We conduct the experiment with Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) on 5000 contexts randomly selected from Government Report (Huang et al., 2021), with full context length ranging from 100–1000 tokens. The ground truth oracle we use here is Prior Work Oracle as introduced in Sec. 5.
@@
To quantitatively assess performance, we compute the maximized Youden’s $J$ statistic (Youden, 1950):
@@
The setting $p=0.90$ achieves the highest Youden’s $J$ ($0.840$), indicating the most balanced trade-off between sensitivity (TPR) and specificity ($1-\text{FPR}$). While $p=0.80$ achieves the highest accuracy (91.9%), its $J$ value is lower, reflecting weaker overall discriminative balance. Therefore, we adopt $p=0.90$ as our standard nucleus sampling configuration in the main experiments. This choice is consistent with prior work recommending $p$ in the 0.9–0.95 range for stable yet diverse generation quality (Holtzman et al., 2020).
@@ F.1 Probability Shift as Selection Metric
A comparison between this form of metric and our probability difference LSPS becomes important when you consider works such as Fang et al. (2025) use the log probability ratio to detect long-short context/next-tokens. Additionally, in Malkin et al. (2022) and Duh et al. (2024) the authors use the log ratio for probability and logits with in the form of $\mathbf{p}(\mathbf{s})_{t}^{1.5}\mathbf{p}_{[-32:]}(\mathbf{s})_{t}^{0.5}(t)$ to improve the final probability of long-context tokens. Implicitly, such methods select and increase the probability of tokens who’s probability ratio increases rather than the difference itself.
In order to compare the choice of metric, we compare the use of LSPS and LSPR when boosting the final word/token of on the validation set of LAMBADA (Paperno et al., 2016) dataset. For each context, we have a ground truth next token, which we call the target token $\hat{t}$. All other tokens we call incorrect tokens. We analyze the choice of which token’s are selected for probability increase rather than the improvement itself, therefore we don’t include multiplicative factors like $\lambda$. For each case of boosting, we consider the following scenarios:
@@
On the other hand, for the probability difference case, we have implicitly set a minimal lower bound on the $\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s})_{t}\geq\epsilon$ by only selecting tokens where $|\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s})_{t}-\operatorname{\mathbf{p}}_{p=0.9}(\mathbf{s}_{[-32:]})_{t}|\geq\epsilon$. This allows us to circumvent tokens who have very low probability where the log ratio behavior could be noisy and misleading. We believe this is a potential reason for why the long context token selection algorithm in Fang et al. (2025) uses $\log\left[\mathbf{p}(\mathbf{s})\right]_{t}\geq 10^{-2}$ ( Long-Context Likelihood (LCL) ) in addition to their log probability ratio to classify tokens.
@@ F.4 Impact of Decoding
F.5 Analysis on additional datasets/models
We provide additional MCL results on more recent models, including Qwen-3-8B Base (Yang et al., 2025) and Mistral-v0.3-7B (Mistral AI, 2024), evaluated on both short- and long-context datasets. These results, shown in Fig. 34, confirm that short-context dominance persists even in these updated model families. We further evaluate TaBoo on two additional QA datasets from L-Eval (An et al., 2023), namely NaturalQuestions, MultiDocQA and ScientificQA, using the same models. To ensure that the QA tasks genuinely require contextual information (i.e., cannot be answered in a closed-book setting), we first run each model on the question alone and remove any example for which the model achieves an F1 score of $\geq 0.4$. As shown in Tab. 10, both models find these QA tasks challenging due to their longer contexts and higher difficulty. Nonetheless, TaBoo consistently improves performance over vanilla inference and CAD, demonstrating that our boosting approach remains effective even on these stronger, more context-heavy benchmarks.